# Tailoring Personality Traits in Large Language Models via Unsupervisedly-Built Personalized Lexicons

### v2

@@ 3.2 Manipulating LLM’s personality
We employ UBPL to manipulate personality during the decoding phase of LLMs. Let $D$ represent the output of the last mapping layer of LLMs. The normalization function (i.e., $Softmaxfunction$) is denoted as $Norm(\cdot)$, the cumulative probability function is denoted as $P(\cdot)$, and $s$ represents the predicted probability of subwords in the vocabulary. $P_{0}$ and $T_{0}$ are model-defined parameters.
In the first step of Top-p nucleus sampling, we obtain the initial candidate word probability vector:
$R_{1}=\text{Norm}(D)=\{s^{1},s^{2},...,s^{|V|}\}$ (10)
Where $Norm(x)$ = $Softmax(x/T_{0})$ , $s^{z}$ represents the probability of subword $w^{z}$ ($z\leq|V|$). Then, this strategy filters out (in reverse order) candidate subwords whose cumulative probability exceeds $P_{0}$, thereby narrowing the sampling space. We express this process using the $f(\cdot)$ function:
$f(R_{1})=\{\max\{P_{0}-P(s_{z}),0\}\cdot\frac{s_{z}}{P_{0}-P(s_{z})}\}_{z=1}^{|V|}$ (11)
Next, we alter the probability vector $R_{1}$ with UBPL, resulting in the final predicted probability vector $(R_{2})$ for the next word with injected personality. This process can be represented by the following mapping:
$f(R_{1})\xrightarrow{G(\cdot)}R_{2}$ (12)
$R_{2}=\{s_{z}\cdot(1+G[L_{\text{val}}(s_{z})])\}_{z=1}^{|V|}$ (13)
@@
$G[L_{val}(s_{z})]=\alpha\cdot\sum_{t=1}^{5}\beta_{t}\cdot L_{val}^{t}(s_{z})$ (14)
After obtaining a new probability vector $R_{2}$ for the next candidate word injected with personality, the next word $W$ is obtained using polynomial sampling from $R_{2}$.
Users can control the overall degree of personality injection through the parameter $\alpha$. Additionally, they have the flexibility to adjust the manifestation of five personality traits exhibited by the model in a finer granularity by manipulating the parameters $\beta_{1}$ through $\beta_{5}$. When $\beta_{t}>0$, it amplifies the expression of trait $t$; conversely, when $\beta_{t}\leq 0$, it diminishes the intensity of trait $t$.
Input: $A$, $sp(\cdot)$, $V$, $S$, $M$, $\alpha$, $D$, $G(\cdot)$, $P_{0}$
Output: $L$, $W$
1Part I: Building UBPL Unsupervisedly $L_{key}\leftarrow V;L_{val}\leftarrow[0,0,0,0,0]^{|V|}$ ;
2 for $i\leftarrow 0$ to $9$ do
3       $j\leftarrow 0$;
4       while $j<200$ do
5             $\{w_{1},w_{2},\ldots\}\leftarrow sp(A_{ij})$;
6             $t=i/2$;
7             foreach $w$ in $\{w_{1},w_{2},\ldots\}$ do
8                   if $i\bmod 2=0$ then
9                         $L[w][t]\leftarrow L[w][t]+1$;
10
11                  else
12                         $L[w][t]\leftarrow L[w][t]-1$;
13
14                   end if
15
16             end foreach
17            $j\leftarrow j+1$;
18
19       end while
20
21 end for
[…]
23       $L_{val}\leftarrow\left\{S\cdot\tanh\left(\frac{v_{z}}{N}\right)\right\}^{|V|}_{z=1}$;
24       $Update$ $N$ $using$ $the$ $Binary$ $Search$;
25
26 end while
27Return: $L$ Part II: Manipulating LLM’s personality $R_{1}\leftarrow\text{Norm}(D)\leftarrow\{s^{1},s^{2},...,s^{|V|}\}$ ;
28 $f(R1)\leftarrow\{\max\{P_{0}-P(s_{z}),0\}\cdot\frac{s_{z}}{P_{0}-P(s_{z})}\}_{z=1}^{|V|}$ ;
29 $R_{2}\leftarrow\{s_{z}\cdot(1+G[L_{\text{val}}(s_{z})])\}_{z=1}^{|V|}$ ;
30 $Sample$ $W$ $from$ $R_{2}$ ;
@@
Algorithm 1 the UBPL method
  Model $-1.0$ $-0.5$ $0.0$ $0.5$ $1.0$ $R$ $P$
Llama2-7b-chat $4.286(0.31)$ $4.343(0.31)$ $4.427(0.31)$ $4.525(0.28)$ $4.558(0.26)$ $0.991$ $031E-$ 03
OpenChat3.5-7b $3.626(0.64)$ $3.756(0.61)$ $3.981(0.44)$ $4.182(0.37)$ $4.237(0.39)$ $0.986$ $032E-$ 03
Neural-chat-7b $3.809(0.58)$ $3.876(0.56)$ $3.999(0.50)$ $4.161(0.44)$ $4.220(0.41)$ $0.989$ $031E-$ 03
Baichuan2-7B-Chat $3.584(0.27)$ $3.710(0.26)$ $4.036(0.38)$ $4.248(0.39)$ $4.336(0.42)$ $0.983$ $033E-$ 03
Llama2-13b-chat $3.856(0.57)$ $3.891(0.54)$ $4.135(0.46)$ $4.298(0.41)$ $4.322(0.38)$ $0.964$ $038E-$ 03
Yi-34b-Chat $4.141(0.42)$ $4.172(0.42)$ $4.246(0.49)$ $4.431(0.40)$ $4.424(0.38)$ $0.941$ $033E-$ 03
Table 1: Single trait manipulating. This table presents the outcomes of single-trait regulation across six models using UBPL. Specifically, it delineates the variations in the mean scores (standard deviations) of five personality traits for the six models as $\alpha=1$, and $-1\leq\beta_{t}\leq 1$. Furthermore, we display the Pearson correlation coefficients (R) and corresponding confidence levels (P) between the mean scores and $\beta_{t}$. Notably, all R values exceed $0.9$, and all P values are below $0.05$, indicating the statistically significant strong correlation between personality trait expression intensity and $\beta_{t}$. This substantiates the effectiveness of our UBPL method in achieving fine-grained control over the expression levels of personality traits in LLMs.
@@ 3.3 Comparison with prior work
As discussed in Section 2.2, there have been two primary methods previously employed to alter the personality of LLMs: fine-tuning and prompt engineering.
In contrast to the fine-tuning, our method obviates the need for resource-intensive parameter fine-tuning. Unlike the inefficiencies inherent in the fine-tuning paradigm, which stem from the requirement to execute fine-tuning steps for each model, UBPL can be seamlessly applied to any open-source LLM in a modular, plug-and-play fashion. In comparison to prompt engineering, UBPL doesn’t necessitate the meticulous design of prompts to coax the model into exhibiting varying degrees of personality traits. Users only need to set $\alpha$ and $\beta$ parameters to regulate the expression intensity of different personality traits in the model at a finer granularity.
The above advantages over the previous methods are based on the effectiveness of our method, so the goal of our experiment is to comprehensively and in detail demonstrate the significant effectiveness of UBPL.
@@ 4.1 LLMs for experiments
To thoroughly demonstrate the effectiveness and generalizability of our method, we conducted experiments on six representative LLMs with model parameters ranging from 7 billion to 34 billion: Llama2-7b-chat(Touvron et al., 2023), OpenChat3.5-7b(Wang et al., 2023), Neural-chat-7b(Intel, 2023), Baichuan2-7B-Chat(Baichuan, 2023), Llama2-13b-chat(Touvron et al., 2023), and Yi-34b-Chat(01.AI, 2023). All the LLMs employ a Top-p nucleus sampling decoding strategy, with a probability threshold ($P_{0}$) of 0.95 and a temperature ($T_{0}$) of 0.85.
@@ 4.2.1 Automatic assessment
To ensure the intrinsic consistency and effectiveness of the assessment, we engaged Llama2-13b-chat in the automatic assessment process. Specifically, we embedded each question of SJTs4LLMs and the corresponding answers generated by the model into Template-2 and asked Llama-13b-chat to score the different personality levels displayed by the model, and finally gathered scores into a five-dimension Likert scale for statistical analysis. Details of Template-2 are in Appendix B.
@@ 4.2.2 Human assessment
Constrained by manpower costs, we recruited a limited cohort of 10 highly educated volunteers for the human assessment process of the Llama7/13b models. At baseline, we randomly selected 40 question-answer pairs for each personality trait theme (constituting 40% of the total) and solicited degree-of-trait ratings from the volunteers. The results were recorded on a five-dimensional Likert scale, and subsequent statistical analysis involved computing the mean and variance.
@@ 5.1 Main results
To demonstrate the effectiveness of our method, we conducted comprehensive experiments. Section 5.1.1 provides a summary of the results demonstrating the manipulation of personality traits across six LLMs using UBPL. In Section 5.1.2, we present the outcomes of UBPL in jointly manipulating the expression of multiple personality traits. In Section 5.1.3, we compare automatic assessment with human assessment to demonstrate the effectiveness of our assessment methodology.
@@ 5.1.1 Single trait manipulating
The results in Table 1 demonstrate the effectiveness of using UBPL to manipulate a single personality of LLMs. Here, $\alpha$ is set to 1, and $|\beta_{t}|\leq 1$ ($\beta_{\neq t}=0$) for $t\in\{OPE,CON,EXT,AGR,NEU\}$. The Pearson correlation coefficients ($R$) are consistently greater than 0.9, signifying a robust positive correlation between $\beta_{t}$ and the intensity of personality expression in LLMs. The confidence level ($P$) is significantly below 0.05, providing compelling evidence that our UBPL can effectively manipulate the intensity of fine-grained personality expression in LLMs. For detailed results on the manipulation of single personality traits for these six LLMs, please see Figure 4.
Figure 4: Detailed results of manipulation of single personality trait. In the “Mean” column, cooler tones indicate smaller values, while warmer tones signify larger values. The table reveals the following observations: 1) Different LLMs exhibit distinct personalities, aligning with previous research findings; 2) When employing our UBPL method, the intensity scores of LLM personalities show a strong positive correlation with the user-controllable $\beta$. This indicates that our UBPL method effectively allows for fine-grained control over the intensity of personality traits expressed by LLMs.
@@ 5.1.2 Multiple trait manipulating
Figure 5: Multiple trait manipulating. The figure above shows the effects of UBPL on multiple personality combinations. In this set of experiments, $\alpha$ was set to 1, and $|\beta_{t}|$ was set to 1. The color tones in the figure represent the expected outcomes based on the personality trait correlations outlined in Table 2, where cool tones indicate that the scores should decrease and warm tones indicate that the scores should increase. The numerical values in the figure depict the changes in the model’s scores on different personality traits compared to the baseline scores after applying the UBPL method. It can be observed that the numerical changes align with the color tones, indicating consistency with the expected results. This demonstrates the effectiveness of our method in the regulation of multiple personalities.
The psychological research results presented in Table 2 reveal interdependencies among the five personality traits within the Big Five personality theory. Consequently, manipulating multiple personality traits is more intricate compared to that of a single personality trait. When we increase the intensity of expression of a specific personality trait, the intensity of expression of other personality traits is also affected.
Considering the adjusted Spearman correlation coefficients ($\rho$) in Table 2, indicating positive correlations among OPE, CON, EXT, and AGR, and negative correlations with NEU, we designed three sets of sub-experiments using Llama2-13b-chat as the model:
@@ 1. Dual Traits Manipulation:
[…]
@@ 2. Triple Traits Manipulation:
[…]
@@ 3. Quadruple Traits Manipulation:
[…]
Why adopt the aforementioned experimental design? Why not manipulate any combination of personality traits and observe the results?
Certainly, users have the flexibility to manipulate any combination of different personality traits of the model at will. However, it is crucial to reiterate that the purpose of our experiment is to demonstrate the effectiveness of UBPL. The evidence in Table 2 demonstrates mutual influences among the five personality traits, such as the strong positive correlation between OPE and EXT. When we set $\beta_{t}$ to increase the strength of OPE and decrease the strength of EXT, regardless of the outcome, we cannot conclusively attribute the results to the impact of UBPL. This is because we have not yet been able to precisely quantify the inter-correlations between personality traits. Therefore, in this context, we collectively enhance or diminish the expression intensity of positively correlated personality traits. This setup ensures that the results can be solely attributed to the effect of the UBPL method, thereby validating its effectiveness.
The experimental results in Table 5 align with the theoretical expectations, affirming the effectiveness of UBPL for the multiple personality manipulating of LLMs.
5.1.3 Human assessment
Figure 6: Comparison of automatic and human assessment. Solid lines show the mean scores of the human assessment, the filled area shows the standard deviation, and the dashed lines show the mean scores of the automatic assessment using LLMs. The results of the automatic assessment and the human assessment are closely aligned, demonstrating the effectiveness of the automatic assessment (Each participant in the human assessment sampled 40 question-answer pairs per trait for each model).
We utilized the Llama2-13b-chat for the automatic assessment of model answers. To demonstrate the effectiveness of this assessment method, we engaged 10 highly qualified individuals in human assessment. Specifically, the human assessment was conducted on models of three different sizes: OpenChat3.5-7b, Llama2-13b-chat, and Yi-34b-Chat. The assessment focused only on the intensity of personality expression in models without UBPL participation.
The comparative results between automatic and human assessment are presented in Figure 6. It is evident from the table that the personality scores obtained through automatic assessment closely align with human assessment results. This substantiates the efficacy of employing LLMs for automatic assessment.

### v3

@@ 3.2 Manipulating LLM’s personality
We employ UPL to manipulate personalities during the decoding phase of LLMs. Let $D$ represent the output of the last mapping layer of the LLMs. The normalization function (i.e., the softmax function) is denoted as $Norm(\cdot)$, the cumulative probability function is denoted as $P(\cdot)$, and $p$ represents the predicted probability of subwords in the vocabulary. $P_{0}$ and $T_{0}$ are model-defined parameters.
In the first step of nucleus sampling, we obtain the initial candidate word probability vector:
$R_{1}=Norm(D)=\{p^{1},p^{2},...,p^{|V|}\}$ (10)
Where $Norm(x)$ = $Softmax(x/T_{0})$ , $p^{z}$ represents the probability of subword $w^{z}$ ($z\leq|V|$). This strategy filters out (in reverse order) candidate subwords whose cumulative probability exceeds $P_{0}$, thereby narrowing the sampling space. We express this process with the $f(\cdot)$ function:
$f(R_{1})=\{\max\{P_{0}-P(w^{z}),0\}\cdot\frac{p_{z}}{P_{0}-P(w^{z})}\}_{z=1}^{|V|}$ (11)
We denote $R_{1}^{\prime}$ as the non-zero part of $f(R_{1})$. Next, we adjust the probability vector $R_{1}^{\prime}$ with UPL, resulting in the final predicted probability vector $R_{2}$ for the next word with injected personality. This mapping is represented as follows:
$R_{1}^{\prime}\xrightarrow{G(\cdot)}R_{2}$ (12)
Specifically, additional probability terms representing personality influence are added:
$R_{2}=\{p_{z}\cdot(1+G(L_{[w^{z}]}))\}_{z=1}^{|V|}$ (13)
@@
$G(L_{[w^{z}]})=\alpha\cdot\sum_{t=1}^{5}\beta_{t}\cdot L_{[w^{z}]}^{t}$ (14)
After obtaining a new probability vector $R_{2}$ for the next candidate word injected with personality, the next word $W$ is determined using polynomial sampling from $R_{2}$.
Users can control the overall degree of personality injection through the parameter $\alpha$. Additionally, they have the flexibility to adjust the manifestation of five personality traits exhibited by the model with finer granularity by manipulating the parameters $\beta_{1}$ through $\beta_{5}$. When $\beta_{t}>0$, it amplifies the expression of trait $t$; conversely, when $\beta_{t}\leq 0$, it diminishes the intensity of trait $t$.
Input: AA, $sp(\cdot)$, $V$, $S$, $M$, $\epsilon$, $D$, $G(\cdot)$, $P_{0}$
Output: UPL: $L$; Next word: $W$
Part I: Building UPL Unsupervisedly
$L_{k}\leftarrow V;L_{v}\leftarrow[0,0,0,0,0]^{|V|}$ ;
for $i\leftarrow 0$ to $9$ do
 $j\leftarrow 0$;
 while $j<200$ do
    $\{w_{1},w_{2},\ldots\}\leftarrow sp(A_{ij})$;
    $t=i/2$;
    foreach $w$ in $\{w_{1},w_{2},\ldots\}$ do
       if $i\bmod 2=0$ then
          $L^{t}_{[w]}\leftarrow L^{t}_{[w]}+1$;
       else
          $L^{t}_{[w]}\leftarrow L^{t}_{[w]}-1$;
         end if
      end foreach
    $j\leftarrow j+1$;
   end while
end for
while $\max_{c\in\{+,-\}}\left\{\text{Avg}^{c}(t)-M_{t}\right\}>\epsilon$ do
 $L_{v}\leftarrow\left\{S\cdot\tanh\left(\frac{v_{z}}{N}\right)\right\}^{|V|}_{z=1}$;
 UpdateUpdate $(N,S)$ withwith thethe BinaryBinary SearchSearch;
end while
Return: $L$
Part II: Manipulating LLM’s personality
$R_{1}\leftarrow\text{Norm}(D)\leftarrow\{s^{1},s^{2},...,s^{|V|}\}$ ;
$f(R1)\leftarrow\{\max\{P_{0}-P(s_{z}),0\}\cdot\frac{s_{z}}{P_{0}-P(s_{z})}\}_{z=1}^{|V|}$ ;
$R_{2}\leftarrow\{s_{z}\cdot(1+G[L_{\text{val}}(s_{z})])\}_{z=1}^{|V|}$ ;
SampleSample $W$ fromfrom $R_{2}$ ;
@@
Algorithm 1 the UPL method
  Model 1.0-1.0 0.5-0.5 $0.0$ $0.5$ $1.0$ $R$ $P$
Llama2-7b-chat $4.286(0.31)$ $4.343(0.31)$ $4.427(0.31)$ $4.525(0.28)$ $4.558(0.26)$ $0.991$ 1E031E-03
OpenChat3.5-7b $3.626(0.64)$ $3.756(0.61)$ $3.981(0.44)$ $4.182(0.37)$ $4.237(0.39)$ $0.986$ 2E032E-03
Neural-chat-7b $3.809(0.58)$ $3.876(0.56)$ $3.999(0.50)$ $4.161(0.44)$ $4.220(0.41)$ $0.989$ 1E031E-03
Baichuan2-7B-Chat $3.584(0.27)$ $3.710(0.26)$ $4.036(0.38)$ $4.248(0.39)$ $4.336(0.42)$ $0.983$ 3E033E-03
Llama2-13b-chat $3.856(0.57)$ $3.891(0.54)$ $4.135(0.46)$ $4.298(0.41)$ $4.322(0.38)$ $0.964$ 8E038E-03
Yi-34b-Chat $4.141(0.42)$ $4.177(0.42)$ $4.243(0.49)$ $4.373(0.40)$ $4.441(0.38)$ $0.982$ 3E033E-03
Table 1: Single trait manipulating. This table presents the results of single-trait regulation across 66 models using UPL. Mean scores (standard deviations) of the $5$ personality traits for these models are shown, where $\alpha=1$, and $1\leq\beta_{t}\leq 1$. Furthermore, we display the Pearson correlation coefficients (R) and corresponding confidence levels (P) between the mean scores and $\beta_{t}$. Notably, all R values exceed $0.9$, and all P values are below $0.05$, indicating the statistically significant strong correlation between personality trait expression intensity and $\beta_{t}$. This substantiates the effectiveness of our UPL method in achieving fine-grained control over the expression levels of personality traits in LLMs.
@@ 3.3 Comparison with prior work
As discussed in Section 2.2, there have been two primary methods previously employed to alter the personality of LLMs: fine-tuning and prompt engineering. In contrast to fine-tuning, our method obviates the heavy need for resource-intensive parameter fine-tuning. Unlike the inefficiencies inherent in the fine-tuning paradigm, which stem from the requirement to execute fine-tuning steps for each model, UPL can be seamlessly applied to the target open-source LLM in a modular, plug-and-play fashion. In comparison to prompt engineering, UPL doesn’t necessitate the meticulous design of prompts to coax the model into exhibiting varying degrees of personality traits. Users only need to set $\alpha$ and $\beta$ parameters to regulate the expression intensity of different personality traits in the model at a finer granularity. The above advantages over the previous methods are based on the effectiveness of UPL.
In this vein, our experiment aims to comprehensively and thoroughly demonstrate the significant effectiveness of UPL.
@@ 4.1 LLMs for experiments
To thoroughly demonstrate the effectiveness and generalizability of our method, we conducted experiments on $6$ representative LLMs with model parameters ranging from 7 billion to 34 billion: Llama2-7b-chat (Touvron et al., 2023), OpenChat3.5-7b (Wang et al., 2023a), Neural-chat-7b (Intel, 2023), Baichuan2-7B-Chat (Baichuan, 2023), Llama2-13b-chat (Touvron et al., 2023), and Yi-34b-Chat (01.AI, 2023).
All the LLMs employ a Top-p nucleus sampling strategy, with a probability threshold ($P_{0}$) of $0.95$ and a temperature ($T_{0}$) of $0.85$.
@@ 4.2.1 Automatic assessment
To ensure the intrinsic consistency and effectiveness of the assessment, we engaged Llama2-13b-chat in the automatic assessment process. Specifically, we embedded each question of STD and the corresponding answers generated by the model into Template-2 and asked the Llama2-13b-chat to score the different personality levels displayed by the model, and finally gathered scores into a five-dimension Likert scale for statistical analysis. Details of Template-2 are in Appendix C.
@@ 4.2.2 Human assessment
Constrained by manpower costs, we recruited a limited cohort of $10$ highly educated volunteers for the human assessment process of the Llama2-7b/13b. At baseline, we randomly selected $40$ question-answer pairs for each personality trait theme (constituting 4040% of the total) and solicited degree-of-trait ratings from the volunteers. The results were recorded on a five-dimensional Likert scale, and subsequent statistical analysis involved computing the mean and variance.
@@ 5.1 Main results
Firstly, we summarize the results demonstrating the effective manipulation of a single personality across six LLMs using UPL (§5.1.1). Subsequently, we discuss the results of manipulating multiple personalities with UPL (§5.1.2). Finally, we compare automated assessment with human evaluation to underscore the effectiveness of our assessment methodology (§5.1.3).
@@ 5.1.1 Single trait manipulating
The results in Table 1 demonstrate the effectiveness of using UPL to manipulate a single personality of LLMs. Here, $\alpha$ is set to 1, and $|\beta_{t}|\leq 1$ ($\beta_{\neq t}=0$) for $t\in\{OPE,CON,EXT,AGR,NEU\}$. The Pearson correlation coefficients ($R$) are consistently greater than $0.9$, signifying a robust positive correlation between $\beta_{t}$ and the intensity of personality expression in LLMs. The confidence level ($P$) is significantly below $0.05$, providing compelling evidence that our UPL can effectively manipulate the intensity of fine-grained personality expression in LLMs. For detailed results on the manipulation of single personality traits for these LLMs, please see Figure 4.
Figure 4: Detailed results of manipulation of single personality trait. In the “Mean” column, cooler tones indicate smaller values, while warmer tones signify larger values. The table reveals the following observations: (i) Different LLMs exhibit distinct personalities, aligning with previous research findings; (ii) When employing UPL, the intensity scores of LLM personalities show a strong positive correlation with the user-controllable $\beta$. This indicates that our method effectively allows for fine-grained control over the intensity of personality traits expressed by LLMs.
@@ 5.1.2 Multiple trait manipulating
Figure 5: Multiple trait manipulating. The figure above shows the effects of UPL on multiple personality combinations. In this set of experiments, $\alpha$ was set to 1, and $|\beta_{t}|$ was set to 1. The colors in the figure represent the expected outcomes based on the personality trait correlations outlined in Table 3, where light blue indicates that the scores should decrease and orange indicates that the scores should increase. The numerical values in the figure depict the changes in the model’s scores on different personality traits compared to the baseline scores after applying the UPL method. It can be observed that the numerical changes align with the color tones, indicating consistency with the expected results. This demonstrates the effectiveness of our method in the regulation of multiple personalities.
The psychological research results presented in Table 3 illustrate interdependencies among the five personality traits within the Big Five personality theory. Consequently, manipulating multiple personality traits is more complex compared to manipulating a single trait, as increasing the intensity of one trait affects the expression intensity of others.
Considering the adjusted Spearman correlation coefficients ($\rho$) in Table 3, indicating positive correlations among OPE, CON, EXT, and AGR, and negative correlations with NEU, we designed three sets of sub-experiments using Llama2-13b-chat as target model:
@@ 1. Dual Traits Manipulation:
$(\downarrow\text{OPE},\uparrow\text{NEU}),(\downarrow\text{CON},\uparrow\text{NEU}),$}
$(\downarrow\text{EXT},\uparrow\text{NEU}),(\downarrow\text{AGR},\uparrow\text{NEU})$}.
@@ 2. Triple Traits Manipulation:
$(\downarrow\text{OPE},\downarrow\text{CON},\uparrow\text{NEU}),$}
$(\downarrow\text{EXT},\downarrow\text{AGR},\uparrow\text{NEU})$}.
@@ 3. Quadruple Traits Manipulation:
$(\downarrow\text{OPE},\downarrow\text{CON},\downarrow\text{EXT},\uparrow\text{NEU}),$}
$(\downarrow\text{CON},\downarrow\text{EXT},\downarrow\text{AGR},\uparrow\text{NEU})$}.
Why adopt the above experimental design? Why not manipulate any combination of personality traits and observe the results? Certainly, users have the flexibility to manipulate any combination of different personality traits of the model at will. However, it is crucial to reiterate that the purpose of our experiment is to demonstrate the effectiveness of UPL. The evidence in Table 3 demonstrates mutual influences among the five personality traits, such as the strong positive correlation between OPE and EXT. When we set $\beta_{t}$ to increase the strength of OPE and decrease the strength of EXT, regardless of the outcome, we cannot conclusively attribute the results to the impact of UPL. This is because the interrelationships between $5$ personality traits cannot be precisely quantified even in quantitative psychology research. Therefore, in this context, we collectively enhance or diminish the expression intensity of positively correlated personality traits. This setup ensures that the results can be solely attributed to the effect of the UPL method, thereby validating its effectiveness.
The experimental results in Table 5 align with the theoretical expectations, affirming the effectiveness of UPL for the multiple personality manipulating of LLMs.
5.1.3 Human Evaluation
Figure 6: Comparison of automatic and human assessment. Solid lines show the mean scores of the human assessment, the filled area shows the standard deviation, and the dashed lines show the mean scores of the automatic assessment using LLMs. The results of the automatic assessment and the human assessment are closely aligned, demonstrating the effectiveness of the automatic assessment.
We utilized the Llama2-13b-chat for the automatic assessment of model personality. To demonstrate the effectiveness of this assessment method, we engaged $10$ highly qualified individuals in human assessment for comparison. Specifically, the human assessment was conducted on models of three different sizes: OpenChat3.5-7b, Llama2-13b-chat, and Yi-34b-Chat. The assessment focused solely on the intensity of personality expression in models without UPL involvement. The comparative results between automatic and human assessments are presented in Figure 6. It is evident that the personality scores obtained through automatic assessment closely align with those from human assessment, substantiating the efficacy of employing LLM for automatic assessment.